OpenStudy (anonymous):

Which expression is equivalent to 9x + 54 – 27 + 3y + 6y?

OpenStudy (tkhunny):

1) DO NOT look at the answers. 2) Simplify the expression. 3) Okay, NOW look at the answers.

OpenStudy (anonymous):

What? lol

OpenStudy (tkhunny):

What wasn't clear about my instructions? Focus on #2. Let's see what you get.

OpenStudy (anonymous):

I Dont know how

OpenStudy (tkhunny):

Yes you do. Solve 54 - 27 = ??

OpenStudy (anonymous):

26

OpenStudy (anonymous):

I got you thanks

OpenStudy (tkhunny):

27 actually. Now, what's 3y + 6y = ??

OpenStudy (anonymous):

Which expression will you generate if you combine the like terms in 5x + 6x + y + y + 3y + 4 + 6 + 2? Can you help me with this?

OpenStudy (tkhunny):

Please identify the like terms. Write them in separate groups.

OpenStudy (anonymous):

9x+27+9y=0